OpenStudy (anonymous):

Find the sum of 5/6 + 1/8 + 3/6 A) 1 17/24 B) 1 7/8 C) 1 1/24 D) 1/2

OpenStudy (shubhamsrg):

and where are you stuck ?

OpenStudy (anonymous):

maybe c i think!

OpenStudy (anonymous):

\[\frac{5}{6}+\frac{1}{8}+\frac{3}{6}\text{= }1+\frac{11}{24} \]

OpenStudy (anonymous):

i kneww it!

OpenStudy (anonymous):

It is not C. C. is 1+1/24, "1 and 1/24th" Mabey a typo?